STEP Data Sheets

/fab/ref/data/test.toc

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1995 T. Toffoli DRAFT, February 2, 1995

Contents

Components (by manufacturer and part number) Fujitsu

  FCN-234P096-G/0 (SBus connector)

2

Fujitsu  FCN-234P096-G/0 [1]

[2] FCN-234P096-G/0  Fujitsu

Fujitsu  Contact Arrangements [1]

[2] Contact Arrangements  Fujitsu

Fujitsu  Panel Cutout Specs [1]

[2] Panel Cutout Specs  Fujitsu